The sum of two numbers is 14. The sum of the squares of the two numbers is 106. Find the two numbers.
Answer by ewatrrr(24785)


Hi
The sum of two numbers is 14
Let x and (14-x) represent the two numbers
Question states***
x^2 + (14-x)^2 = 106
solving for x
2x^2 -28x + 196 = 106
2x^2 -28x + 90 = 0
2(x^2 -14x + 45) = 0
factoring
(x-9)(x-5) = 0 Numbers are 5 and 9
CHECKING our Answer***
25 + 81 = 106